OpenStudy (anonymous):

Add. 18 5/12 + 8 10/17 + 23 7/12 + 28 7/17 A. 77 B. 77 1/2 C. 78 1/2 D. 79

OpenStudy (anonymous):

18(5/12) = (18*12)+5)/12 use this method to make all the numbers in p/q form then take least common multiple of the denominator and then you can solve the question

OpenStudy (anonymous):

Thank you sooo much i have been stuck on a question all day thank you

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

wait hold on the answer you gave me is not on the sheet. let me edit the question

OpenStudy (anonymous):

there now answer it

OpenStudy (anonymous):

\[18(5/12) =((18*12)+5)/12\] = 221/12 \[8(10/17) = ((8*17) +10)/17 = 146/17\] \[23(7/12)= ((23*12) +7)/12 = 283/12\] \[28(7/17) = ((28*17)+7)/17 = 483/17\] (221/12) + (283/12) = (504/12) = 42 (146/17) + (483/17) = (629/17) = 37 42+37 = 79